# Dual process in the two-parameter Poisson–Dirichlet diffusion

### v5

@@
( January 10, 2024)
The two-parameter Poisson–Dirichlet diffusion takes values in the infinite ordered simplex and extends the celebrated infinitely-many-neutral-alleles model, with a two-parameter Poisson–Dirichlet stationary distribution. Here we identify a dual process for this diffusion and obtain its transition probabilities. The dual is shown to be given by Kingman’s coalescent with mutation, conditional on a given configuration of leaves. Interestingly, the dual does not depend on the additional parameter of the stationary distribution. After discussing the sampling probabilities of a two-parameter Poisson–Dirichlet partition drawn conditionally on another partition, we use these together with the dual process to derive the transition density of the diffusion. Our derivation provides a new probabilistic proof of this result, leveraging on an extension of Pitman’s Pólya urn scheme whereby the urn is split after a finite sequence and two urns are run independently onwards. The proof strategy exemplifies the power of duality and could be exported to other models where a dual is available.
@@ 1 Introduction
The two-parameter Poisson–Dirichlet diffusion (later simply two-parameter diffusion), was introduced by [61] and extends the celebrated infinitely-many-neutral-alleles model of [20] (here obtained when $\alpha=0$) to the case of two parameters $\alpha\in[0,1)$ and $\theta>-\alpha$. This diffusion takes values in the closure of the infinite-dimensional ordered simplex taken with respect to the product topology of $[0,1]^{\infty}$, namely
@@
The domain of $L_{\alpha,\theta}$ is taken to be the algebra generated by constants and functions $\varphi_{k}(x)=\sum_{i=1}^{\infty}x_{i}^{k}$ for $k\geq 2$, which is a dense subalgebra of $C(\overline{\nabla}_{\infty})$ and thus forms a core for the closure of $L_{\alpha,\theta}$.
@@
for every Borel set $A$ of $S$, where $k$ is the number of distinct values $y_{j}^{*}$ in $y^{(n)}$, with $j=1,\ldots,k$, with respective observed multiplicities $n_{j}$. The scheme reduces to the original Blackwell–MacQueen generalized Pólya urn scheme when $\alpha=0$. The distribution $\mathrm{PD}_{\alpha,\theta}$ and its corresponding urn scheme (2) have found numerous applications in a variety of fields. We refer the reader to the monographs, and references therein, [7] for fragmentation and coalescent theory, [64] for excursion theory and combinatorics, [55] for Bayesian inference, [70] for machine learning, [24] for stochastic population dynamics. Probably owing to the popularity of the two-parameter Poisson–Dirichlet distribution, since the contribution of [61] there has been a continuous and renewed interest in understanding the finer properties of the two-parameter diffusion, together with other closely related dynamical structures. See [67, 27, 18, 12, 25, 28, 34, 31, 66, 32, 33]. Here it is important to note that the infinitely-many-neutral-alleles model, corresponding to (1) with $\alpha=0$, can be seen as the unlabeled version of a Fleming–Viot diffusion, where the labels that represent the information on individual values in the underlying evolving population are lost, and only the relative frequencies of types (arranged in decreasing order) are retained. Cf. [22], Theorem 9.2.1, or [24], Theorem 5.6. A similar correspondence for the two-parameter diffusion is still object of investigation today, and while [34, 31, 32, 33] have recently provided much insight, many aspects are still to be fully understood.
@@
holds for all $x\in E,d\in F$ and $t\geq 0$. See [49] for a review. In the above identity, the expectation on the left hand side is taken with respect to the law of $X_{t}$, conditional on $X_{0}=x$, while that on the right hand side it is taken with respect to law of $D_{t}$, conditional on $D_{0}=d$. The function $h(x,d)$ satisfying the identity is referred to as the duality function between $X$ and $D$. Duality is an important tool in the theory of stochastic processes [21, 14], which has found widespread applications, among other areas, in mathematical population genetics [56, 5, 2, 48, 15, 8, 16, 26, 10, 23, 11], in statistical physics and interacting particle systems [36, 37, 38, 57, 47, 35], and in statistical inference [58, 59, 3, 53, 4, 54]. Knowledge of the dual process and of its properties therefore has key implications, which range from the study of the well-posedness of the martingale problem associated to $X_{t}$, to the possibility of conducting sequential inference on $X_{t}$ given a finite computational budget, to the ability of deriving the transition function of $X_{t}$ through limit arguments. The latter will also be the main application of our results.
@@
Here we relate the dual process of (1) to Kingman’s coalescent [52]. The literature on Kingman’s coalescent and its extensions is vast, and here we simply refer the reader to the reviews [6, 43] and references therein. As shown later (cf. Theorem 3.3) the dual process to the two-parameter diffusion turns out to coincide with the process that counts, in Kingman’s coalescent, the number of non-mutant lineages ancestral to a sample from the current generation, conditionally given the sample’s unlabelled allelic partition. In particular, the dual process identified here does not depend on the additional parameter $\alpha$. This fact may somewhat be surprising. Indeed, it is well-known that a Kingman’s coalescent tree with $n$ leaves can be realised by sampling lineages sequentially, starting from the root, according to the Blackwell–MacQueen Pólya urn scheme, i.e., (2) with $\alpha=0$. Given the connection between (2) and the stationary distribution of the two-parameter diffusion, it would thus be natural to expect that $\alpha$ plays a role in the law of the dual process. In fact, this parameter has been shown to have a key role in the construction of the diffusion [67, 12] or of its labelled counterpart [34]. A possible explanation is related to the fact that the dual process describes the conditional genealogy of a sample given its allelic partition. As an effect of exchangeability, conditionally on the frequencies, the parameters only contribute to the law of the coalescent’s block-counting process, keeping track of the total number of surviving lineages, regardless of their types. Our results then indicate that the positive and negative effects of $\alpha$ on the block-counting rates balance each other out perfectly giving a zero net effect.
@@
The paper is organized as follows. In Section 2 we collect some notation together with some necessary preliminary notions. In Section 3 we fully characterize the dual process of the two-parameter Poisson–Dirichlet diffusion. In Section 4 we discuss the sampling distributions related to drawing a partition from the two-parameter Poisson–Dirichlet model conditional on another, already observed, partition. These distributions are then used in Section 5 where we derive the transition density of the two-parameter diffusion using the identified dual. This transition density was found in [29] through analytical tools. Together with the dual process, our derivation is instead via a probabilistic approach, which exploits a “split”version of Pitman’s two-parameter urn scheme, that is a system of two such urns, conditionally independent given the same vector of initial observations. The split urns give, in the limit, a bivariate distribution with identical $\mathrm{PD}_{\alpha,\theta}$ marginal laws, whereby the size of the common initial sample measures the strength of the dependence between the two coordinates. It turns out that the transition density of the two-parameter diffusion can be interpreted as the conditional distribution of one coordinate given the other in such a construction, where the size of the initial sample is random and its distribution is given by the block-counting process of the dual. Our probabilistic derivation of the transition density therefore exemplifies the power of duality and could be reproduced, at least in principle, to identify the transition density for models such that this is unknown, but a dual is nonetheless available, as for example is the case of coupled Wright–Fisher diffusions [23].
@@ 2 Preliminaries
A vector $\eta=(\eta_{1},\cdots,\eta_{d})\in\mathbb{N}^{d},d\in\mathbb{Z}_{+}$, is called an integer partition of $n\in\mathbb{Z}_{+}$ if $\eta_{1}\geq\cdots\geq\eta_{d}>0$ and $|\eta|:=\sum_{i=1}^{d}\eta_{i}=n$. When useful to make explicit, the length of $\eta$ will also be denoted as $l(\eta)=d$. Define $\Gamma:=\cup_{n\geq 0}\Gamma_{n}$, where $\Gamma_{0}=\{\emptyset\}$ and $\Gamma_{n}$ is the set of all integer partitions of $n$. It will be convenient to view a partition $\eta$ as an infinite vector obtained by appending an infinite sequence of zero coordinates to the $l(\eta)$-th part of $\eta$. For $\omega,\eta\in\Gamma$, we say that $\omega\subset\eta$ if and only if $\omega_{i}\leq\eta_{i}$ for all $i\geq 1$, so that $(\Gamma,\subset)$ becomes a partially ordered set. Throughout the paper, we will have $n=|\eta|$ and $w=|\omega|$. Given $\eta\in\Gamma$ and $x\in\nabla_{\infty}:=\{x\in\overline{\nabla}_{\infty}:\sum_{i\geq 1}x_{i}=1\}$, define now
@@
and let
@@
where $a_{k}(\eta)$ is the number of groups in $\eta$ with size $k$, so that $\sum_{k=1}^{n}a_{k}(\eta)=d$ and $\sum_{k=1}^{n}ka_{k}(\eta)=n$. Note that for every $x$, $\vec{P}_{\eta}(x)$ is a symmetric function in $\eta$, and for every $\eta$, $\vec{P}_{\eta}(x)$ is a symmetric function in $x$.
Note also that we can extend both $P_{\eta}(x)$ and $\vec{P}_{\eta}(x)$ continuously from $\nabla_{\infty}$ to $\overline{\nabla}_{\infty}$, since $\nabla_{\infty}$ is dense in $\overline{\nabla}_{\infty}$. Cf. [61]. With a slight abuse of notation, we will use the same symbols to denote their continuous extensions. Here $x\in\nabla_{\infty}$ may be regarded as the vector of relative frequencies in an infinite population with potentially infinitely-many types, or of color frequencies for balls in an urn. If we sample $n$ balls with replacement from the urn, then $\vec{P}_{\eta}(x)$ is the probability of observing the color configuration $\eta$ in the sample.
A partition structure is a family of distributions $\{M_{n}(\eta),\eta\in\Gamma_{n}\}$ for $n\geq 1$, that satisfies the consistency condition whereby, for $\omega\in\Gamma_{n-1}$,
(6) $M_{n-1}(\omega)=\sum_{\eta\in\Gamma_{n}:\eta\supset\omega}\frac{\binom{n-1}{\omega}\chi(\omega,\eta)}{\binom{n}{\eta}}M_{n}(\eta)$
where $\chi(\omega,\eta)$ equals $a_{\eta_{i}}(\eta)$ if $\eta_{i}-\omega_{i}=1$ and $\eta_{j}-\omega_{j}=0$ for $j\neq i$, and zero elsewhere. A similar consistency property holds for functions $\eta\mapsto\vec{P}_{\eta}(x)$, given $x\in\nabla_{\infty}$, as well, which are also partition structures. By Kolmogorov’s consistency condition, we can establish a probability measure $\mathcal{M}(\cdot)$ on integer partitions $\Gamma$. We say that a $\Gamma$-valued stochastic process $\{D_{n},n\geq 1\}$ has distribution $\mathcal{M}$, if $D_{n}\in\Gamma_{n}$ and has marginal distribution $M_{n}(\eta)$.
@@
where $\mu$ is a probability measure on $\overline{\nabla}_{\infty}$, called representing measure. Furthermore, if $\{D_{n},n\geq 1\}$ is a family of random partitions with distributions $\{M_{n},n\geq 1\}$, then $Z:=\lim_{n\to\infty}D_{n}/n$ exists almost surely and has distribution $\mu$.
We record the following Lemma which will be used later.
@@
Proof.
Let $D_{n}$ be a family of $\Gamma$-valued random variables with distribution $M_{n}$. Then by Kingman’s representation theorem, we know $D_{n}/n$ converges to $Z$ almost surely and the distribution of $Z$ is $\mu$. Therefore, for any bounded continuous function $f$ on the Kingman simplex $\overline{\nabla}_{\infty}$, we have
$\lim_{n\to\infty}\int_{\overline{\nabla}_{\infty}}f(x)\mu_{n}(dx)=\lim_{n\to\infty}\mathbb{E}\Big{[}f(D_{n}/n)\Big{]}=\mathbb{E}\big{[}f(Z)\big{]}=\int_{\overline{\nabla}_{\infty}}f(x)\mu(dx),$
yielding the result. ∎
When partitions are generated by two-parameter Poisson–Dirichlet distributions, (7) holds with $\mu$ given by $\mathrm{PD}_{\alpha,\theta}$, and $M_{n}(\eta)$ is the probability of observing a partition $\eta$ of $n$ from a sample of size $n$ drawn from the two-parameter Chinese restaurant process. This is the law of the unlabelled empirical frequencies induced by the generalized Pólya urn scheme (2), and yields the Ewens–Pitman sampling formula
@@
where $a_{(n)}=a(a+1)\cdots(a+n-1)$, $a_{(0)}=1$. The expectation of $P_{\eta}$ with respect to $\mathrm{PD}_{\alpha,\theta}$ yields instead the so-called exchangeable partition probability function
@@
Cf. (36) and (42) in [63]. Later we will use the shorter notation $\mathbb{E}_{\alpha,\theta}[{P}_{\eta}]$ and $\mathbb{E}_{\mu}[\vec{P}_{\eta}]$.

### v6

@@
( October 1, 2024)
The two-parameter Poisson–Dirichlet diffusion takes values in the infinite ordered simplex and extends the celebrated infinitely-many-neutral-alleles model, having a two-parameter Poisson–Dirichlet stationary distribution. Here we identify a dual process for this diffusion and obtain its transition probabilities. The dual is shown to be given by Kingman’s coalescent with mutation, conditional on a given configuration of leaves. Interestingly, the dual depends on the additional parameter of the stationary distribution only through the test functions and not through the transition rates. After discussing the sampling probabilities of a two-parameter Poisson–Dirichlet partition drawn conditionally on another partition, we use these notions together with the dual process to derive the transition density of the diffusion. Our derivation provides a new probabilistic proof of this result, leveraging on an extension of Pitman’s Pólya urn scheme, whereby the urn is split after a finite number of steps and two urns are run independently onwards. The proof strategy exemplifies the power of duality and could be exported to other models where a dual is available.
@@ 1 Introduction
The two-parameter Poisson–Dirichlet diffusion (later simply two-parameter diffusion), was introduced by [60] and extends the celebrated infinitely-many-neutral-alleles model of [20] (here obtained when $\alpha=0$) to the case of two parameters $\alpha\in[0,1)$ and $\theta>-\alpha$. This diffusion takes values in the closure of the infinite-dimensional ordered simplex, taken with respect to the product topology of $[0,1]^{\infty}$, namely
@@
The domain of $L_{\alpha,\theta}$ is taken to be the algebra generated by constants and functions $\varphi_{k}(x)=\sum_{i=1}^{\infty}x_{i}^{k}$ for $k\geq 2$. This is a dense subalgebra of $C(\overline{\nabla}_{\infty})$ (cf., e.g., [20], end of page 435), and [60] showed that the closure of $L_{\alpha,\theta}$ in $C(\overline{\nabla}_{\infty})$ generates a Feller semigroup.
@@
for every Borel set $A$ of $S$, where $k$ is the number of distinct values $y_{j}^{*}$ in $y^{(n)}$, with $j=1,\ldots,k$, with respective observed multiplicities $n_{j}$. The scheme reduces to the original Blackwell–MacQueen generalized Pólya urn scheme when $\alpha=0$. The distribution $\mathrm{PD}_{\alpha,\theta}$ and its corresponding urn scheme (2) have found numerous applications in a variety of fields. We refer the reader to the following monographs, and references therein: [7] for fragmentation and coalescent theory, [63] for excursion theory and combinatorics, [54] for Bayesian inference, [68] for machine learning, [24] for stochastic population dynamics. Probably owing to the popularity of the two-parameter Poisson–Dirichlet distribution, since the contribution of [60] there has been a continuous and renewed interest in understanding the finer properties of the two-parameter diffusion, together with other closely related dynamical structures. See [66, 27, 18, 12, 25, 28, 33, 30, 65, 31, 32]. Here it is important to recall that the infinitely-many-neutral-alleles model, corresponding to (1) with $\alpha=0$, can be seen as the unlabeled version of a Fleming–Viot diffusion, where the labels that represent the information on individual values in the underlying evolving population are lost, and only the relative frequencies of types (arranged in decreasing order) are retained. Cf. [22], Theorem 9.2.1, or [24], Theorem 5.6. A similar correspondence for the two-parameter diffusion is still object of investigation today, and while [33, 30, 31, 32] have recently provided advancements, many aspects are still to be fully understood.
@@
holds for all $x\in E,d\in F$ and $t\geq 0$. See [48] for a review. In the above identity, the expectation on the left hand side is taken with respect to the law of $X_{t}$, conditional on $X_{0}=x$, while that on the right hand side is taken with respect to the law of $D_{t}$, conditional on $D_{0}=d$. The function $h(x,d)$ satisfying the identity is referred to as the duality function between $X$ and $D$. Duality is an important tool in the theory of stochastic processes [21, 14], which has found widespread applications, among other areas, in mathematical population genetics [55, 5, 2, 47, 15, 8, 16, 26, 10, 23, 11], in statistical physics and interacting particle systems [35, 36, 37, 56, 46, 34], and in statistical inference [57, 58, 3, 53, 4, 52]. Knowledge of the dual process and of its properties therefore has key implications, which range from the study of the well-posedness of the martingale problem associated to $X_{t}$, to the possibility of conducting sequential inference on $X_{t}$ given a finite computational budget, to the ability of deriving the transition function of $X_{t}$ through limit arguments. The latter will also be the main application of our results.
@@
Here we relate the dual process of (1) to Kingman’s coalescent [51]. The literature on Kingman’s coalescent and its extensions is vast, and here we simply refer the reader to the reviews [6, 42] and references therein. As shown later (cf. Theorem 3.3) the dual process to the two-parameter diffusion turns out to coincide with the process that counts, in Kingman’s coalescent, the number of non-mutant lineages ancestral to a sample from the current generation, conditionally given the sample’s unlabelled allelic partition. In particular, the dual transition rates still depend only on the parameter $\theta$, like for the $\alpha=0$ model, while the dependence on the additional parameter $\alpha$ is through the test (or duality) functions. This fact may somewhat be surprising. Indeed, it is well-known that a Kingman’s coalescent tree with $n$ leaves can be realised by sampling lineages sequentially, starting from the root, according to the Blackwell–MacQueen Pólya urn scheme, i.e., (2) with $\alpha=0$. Given the connection between (2) and the stationary distribution of the two-parameter diffusion, it would thus be natural to expect that $\alpha$ plays a role in the law of the dual process. In fact, this parameter has been shown to have a key role in the construction of the diffusion [66, 12] or of its labelled counterpart [33]. One would then expect $\alpha$ to influence the deletion of groups both through mutation and coalescence. Our results instead indicate that these two effects of $\alpha$ on the block-counting rates balance each other out perfectly, giving a zero net effect. A possible explanation is related to the fact that the dual process describes the conditional genealogy of a sample given its allelic partition. As an effect of exchangeability, conditionally on the frequencies, the parameters only contribute to the law of the coalescent’s block-counting process, keeping track of the total number of surviving lineages, regardless of their types.
@@
The paper is organized as follows. In Section 2 we collect some notation together with some necessary preliminary notions. In Section 3 we fully characterize the dual process of the two-parameter diffusion. In Section 4 we discuss the sampling distributions related to drawing a partition from the two-parameter Poisson–Dirichlet model conditional on another, already observed, partition. These distributions, together with the dual process, are then used in Section 5 where we derive the transition density of the two-parameter diffusion using the identified dual. This transition density was found in [29] through analytical tools. Our derivation uses instead a probabilistic approach, which exploits a “split”version of Pitman’s two-parameter urn scheme (2), namely is a system of two such urns that grow conditionally independently given the same vector of initial observations. The split urns give, in the limit, a bivariate distribution with identical $\mathrm{PD}_{\alpha,\theta}$ marginal laws, whereby the size of the common initial sample measures the strength of the dependence between the two coordinates. It turns out that the transition density of the two-parameter diffusion can be interpreted as the conditional distribution of one urn composition given the other in such a construction, where the size of the initial sample is random and its distribution is given by the block-counting process of the dual. Our probabilistic derivation of the transition density therefore exemplifies the power of duality and could be reproduced, at least in principle, to identify the transition density for models when this is unknown, but a dual is nonetheless available, e.g., for coupled Wright–Fisher diffusions [23].
@@ 2 Preliminaries
A vector $\eta=(\eta_{1},\cdots,\eta_{d})\in\mathbb{N}^{d},d\in\mathbb{Z}_{+}$, is called an integer partition of $n\in\mathbb{Z}_{+}$ if $\eta_{1}\geq\cdots\geq\eta_{d}>0$ and $|\eta|:=\sum_{i=1}^{d}\eta_{i}=n$. When useful to make explicit, the length of $\eta$ will also be denoted as $l(\eta)=d$. Define $\Gamma:=\cup_{n\geq 0}\Gamma_{n}$, where $\Gamma_{0}=\{\emptyset\}$ and $\Gamma_{n}$ is the set of all integer partitions of $n$. It will be convenient to view a partition $\eta$ as an infinite vector obtained by appending an infinite sequence of zero coordinates to the $l(\eta)$-th part of $\eta$. For $\omega,\eta\in\Gamma$, we say that $\omega\subset\eta$ if and only if $\omega_{i}\leq\eta_{i}$ for all $i\geq 1$, so that $(\Gamma,\subset)$ becomes a partially ordered set. Throughout the paper, we will have $n=|\eta|$ and $\ell=|\omega|$. Given $\eta\in\Gamma$ and $x\in\nabla_{\infty}:=\{x\in\overline{\nabla}_{\infty}:\sum_{i\geq 1}x_{i}=1\}$, define now
@@
so that in particular when $\eta$ is given by a singleton we have $P_{(1)}=1$, and for definiteness we also set ${P}_{\emptyset}(x):=1$ (note however that $\eta=\emptyset$ will not be needed; cf. Section 3). Let also
@@
where $a_{k}(\eta)$ is the number of groups in $\eta$ with size $k$, so that $\sum_{k=1}^{n}a_{k}(\eta)=d$ and $\sum_{k=1}^{n}ka_{k}(\eta)=n$. Note that for every $x$, $\vec{P}_{\eta}(x)$ is a symmetric function in $\eta$, and for every $\eta$, it is a symmetric function in $x$. Here and throughout we adopt the same convention used in [20, 60], whereby both functions $P_{\eta}(x)$ and $\vec{P}_{\eta}(x)$ can be extended continuously from $\nabla_{\infty}$ to $\overline{\nabla}_{\infty}$, since $\nabla_{\infty}$ is dense in $\overline{\nabla}_{\infty}$. So for example $P_{(1)}$ equals $1$ and not $\sum_{i\geq 1}x_{i}$, which is not continuous on $\overline{\nabla}_{\infty}$. With a slight abuse of notation, we will use the same symbols to denote their continuous extensions onto $\overline{\nabla}_{\infty}$.
Here $x\in\nabla_{\infty}$ may be regarded as the vector of relative frequencies in an infinite population with potentially infinitely-many types, or of color frequencies for balls in an urn. If we sample with replacement $n$ items from the population or $n$ balls from the urn, given $x$, then $\vec{P}_{\eta}(x)$ is the probability that the sample features $\eta_{1}$ items/balls of one type/color (regardless of which), $\eta_{2}$ items/balls of a different type/color, etc.
These function assume a prominent role in the framework of partition structures. A partition structure is a family of distributions $\{M_{n}(\eta),\eta\in\Gamma_{n}\}$ for $n\geq 1$, that satisfies the consistency condition whereby, for $\omega\in\Gamma_{n-1}$,
(6) $M_{n-1}(\omega)=\sum_{\eta\in\Gamma_{n}:\eta\supset\omega}\frac{\binom{n-1}{\omega}\chi(\omega,\eta)}{\binom{n}{\eta}}M_{n}(\eta),$
where $\chi(\omega,\eta)$ equals $a_{\eta_{i}}(\eta)$ if $\eta_{i}-\omega_{i}=1$ and $\eta_{j}-\omega_{j}=0$ for $j\neq i$, and zero elsewhere. Here $M_{0}(\emptyset)=M_{1}((1))=1$. A similar consistency property holds for functions $\eta\mapsto\vec{P}_{\eta}(x)$, given $x\in\nabla_{\infty}$, as well, which are also partition structures. By Kolmogorov’s consistency condition, we can establish a probability measure $\mathcal{M}(\cdot)$ on integer partitions $\Gamma$. We say that a $\Gamma$-valued stochastic process $\{D_{n},n\geq 1\}$ has distribution $\mathcal{M}$, if $D_{n}\in\Gamma_{n}$ and has marginal distribution $M_{n}(\eta)$.
@@
where $\mu$ is a probability measure on $\overline{\nabla}_{\infty}$, called representing measure. Furthermore, if $\{D_{n},n\geq 1\}$ is a family of random partitions with distributions $\{M_{n},n\geq 1\}$, then $Z:=\lim_{n\to\infty}D_{n}/n$ exists almost surely and has distribution $\mu$. For ease of later reference, we state without proof the following Lemma, which is an immediate consequence of the above.
@@
When partitions are generated by two-parameter Poisson–Dirichlet distributions, (7) holds with $\mu$ given by $\mathrm{PD}_{\alpha,\theta}$, and $M_{n}(\eta)$ is the probability of observing a partition $\eta$ of $n$ from a sample of size $n$ drawn from the two-parameter Chinese restaurant process. This is the law of the unlabelled empirical frequencies induced by the generalized Pólya urn scheme (2), and yields the Ewens–Pitman sampling formula [61]
@@
where $a_{(n)}=a(a+1)\cdots(a+n-1)$, $a_{(0)}=1$. The expectation of $P_{\eta}$ with respect to $\mathrm{PD}_{\alpha,\theta}$ yields instead the so-called exchangeable partition probability function [61]
@@
Cf. also (36) and (42) in [62]. Later we will use the shorter notation $\mathbb{E}_{\alpha,\theta}[{P}_{\eta}]$ and $\mathbb{E}_{\mu}[\vec{P}_{\eta}]$.